1. Solve for x
a) √(x+1) + √(2x-5) = 3
Can you please help me out? Thanks so much in advance:)
Can you please show the steps it would really help me understand:)

Answer by stanbon(75887)

√(x+1) + √(2x-5) = 3
-------------------------
Square both sides to get:
(x+1) + (2x-5) + 2*sqrt[(x+1)(2x-5)] = 9
--------
2sqrt[2x^2 -3x -5] = 13 - 3x
-----
Square both sides again:
4[2x^2-3x-5] = 169-78x+9x^2
------
8x^2 - 12x -20 = 169 - 78x + 9x^2
x^2 - 66x +189 = 0
----
(x-3)(x+63) = 0
----
Positive solution:
x = 3
===============
Check the answer in the original equation.
====================
Cheers,
Stan H.
====================